# supplement

### v1

@@ Abstract
In this paper, we address the challenge of Markov Chain Monte Carlo (MCMC) algorithms within the Approximate Bayesian Computation (ABC) framework, which often get trapped in local optima due to their inherent local exploration mechanism. We propose a novel Global-Local ABC-MCMC algorithm that combines the “exploration” capabilities of global proposals with the “exploitation” finesse of local proposals. By integrating iterative importance resampling into the likelihood-free framework, we establish an effective global proposal distribution. We select the optimum mixture of global and local moves based on a unit cost version of expected squared jumped distance via sequential optimization. Furthermore, we propose two adaptive schemes: The first involves a normalizing flow-based probabilistic distribution learning model to iteratively improve the proposal for importance sampling, and the second focuses on optimizing the efficiency of the local sampler by utilizing Langevin dynamics and common random numbers. We numerically demonstrate that our method improves sampling efficiency and achieve more reliable convergence for complex posteriors. A software package implementing this method is available at https://github.com/caofff/GL-ABC-MCMC.
Keywords: Approximate Bayesian computation, common random numbers, iterative sampling importance resampling, normalizing flow, sequential optimization.
@@ 1 Introduction
Traditional Bayesian inference typically relies on the assumption that the likelihood functions of statistical models are estimable. However, for many complex applications, these likelihood functions either lack explicit expressions or are difficult to estimate numerically. Approximate Bayesian computation (ABC) (Pritchard et al., 1999; Beaumont et al., 2002) is a likelihood-free inference method based on simulations. It only requires simulating synthetic data from the model without evaluating the likelihood function. Rejection ABC (Tavaré et al., 1997; Pritchard et al., 1999) is the simplest and most intuitive version of ABC algorithms, which involves repeatedly drawing parameter samples ${\boldsymbol{\theta}}$ independently from prior $\pi$, simulating synthetic data $\boldsymbol{x}$ for each sample of ${\boldsymbol{\theta}}$, and rejecting the parameter ${\boldsymbol{\theta}}$ if the discrepancy $\mathfrak{D}(\boldsymbol{x},\boldsymbol{y})$ between the observed data $\boldsymbol{y}$ and the simulated data $\boldsymbol{x}$ exceeds a pre-specified tolerance threshold $\varepsilon$. In practice, sampling parameters from a prior distribution can be inefficient, especially in high dimensional cases, as most of samples drawn from the prior distribution fall into the low posterior regions, resulting very high rejection rates. Consequently, accepting one sample may require thousands or even millions of draws.
@@
Marjoram et al. (2003) introduce an ABC Markov chain Monte Carlo (MCMC) algorithm to approximate the posterior distribution. Proposal distributions are typically chosen to generate local moves that depend on the last state of the chain to ensure an admissible acceptance rate. Theoretically, a small tolerance value $\varepsilon$ leads to a good posterior approximation. However, designing an efficient proposal to explore the parameter space, especially in the case of high dimensional multimodal posterior distributions, poses a significant challenge. Wegmann et al. (2009) enhanced the performance of ABC-MCMC by relaxing the tolerance, incorporating subsampling, and applying regression adjustments to the MCMC output. Sisson et al. (2007); Del Moral et al. (2012) integrated ABC-MCMC algorithms into the SMC framework (Del Moral et al., 2006). The sequence of intermediate target distributions is defined by ABC posteriors with a series of tolerance parameters. Del Moral et al. (2012) adaptively determined this sequence by managing the divergence of particles. Clarté et al. (2020); Rodrigues et al. (2020) proposed Gibbs version of the ABC approaches that focuses on lower-dimensional conditional distributions. Hamiltonian Monte Carlo (HMC) (Duane et al., 1987; Neal et al., 2011) employs Hamiltonian dynamics to propose candidate parameters, efficiently avoiding the random walk behavior that can impede the mixing of proposals in high-dimensional spaces. (Meeds et al., 2015) explored the application of HMC to ABC and proposed using completely random numbers (CRNs) to improve the stability of gradient estimates to improve Hamiltonian dynamic.
@@
However, for a high dimensional multimodal posterior distribution, all local samplers mix poorly because the exploration is inherently slow, and switching between isolated modes can be extremely rare. Independent proposals can generate more global updates to transfer particles between modes, but they are difficult to design. A poor designed independent sampler in MCMC results in an inefficient random walk, causing the algorithm to converge slowly. For likelihood-based Bayesian inference, Samsonov et al. (2022) improved an Explore-Exploit MCMC strategy that couples a local move and a global move in each iteration. Gabrié et al. (2022) performed a global move after a fixed number of local moves during MCMC iterations. Moreover, normalizing flows (Rezende and Mohamed, 2015; Papamakarios et al., 2021) are used to improve the exploration capabilities of MCMC independent proposals.
In this paper, we propose a novel approach that combines the advantages of global MCMC proposal and local MCMC proposal for challenging posterior distribution within the likelihood-free framework. First, we introduce a likelihood-free iterative-sampling importance resampling (i-SIR) approach to construct an efficient global kernel for ABC-MCMC. We demonstrates that the ABC-i-SIR algorithm can be interpreted as a systematic-scan two-stage Gibbs sampler and prove the uniform geometric ergodicity of the ABC-i-SIR Markov kernel. Furthermore, we also establish the V-geometric ergodicity of the combination of the ABC-i-SIR Markov kernel and the local Markov kernel, and thus show that under certain conditions, the mixing rate of the Global-Local Markov Kernel is significantly better than that of the local Markov kernel. Second, we propose a sequential optimization approach based on a unit cost version of expected square jump distance (cESJD) criterion to select the hyper-parameters of Global-Local ABC-MCMC (GL-ABC-MCMC), such as the batch size of ABC-i-SIR and the mixture proportion of global and local moves. Third, we develop two adaptive schemes to enhance the efficiency of GL-ABC-MCMC by automatically constructing the local and global proposal distributions. The first scheme improves the importance proposal distribution of ABC-i-SIR using normalizing flows during the iterations of the Markov chain. The second scheme applies the Metropolis-adjusted Langevin algorithm to ABC, and utilizes common random numbers to enhance the estimate of the gradient of the log-likelihood. Our numerical experiments indicate that the proposed method can mix faster than existing methods.
@@
The rest of paper is organized as follows. Section 2 presents our GL-ABC-MCMC, including ABC-i-SIR, the sequential optimization of hyper-parameters and the theoretical properties. Section 3 introduces two adaptive schemes for improving the global and local proposal distributions. Sections 4 and 5 demonstrate the effectiveness of our method through synthetic examples studies and two real model analysis. Section 6 provides the conclusion, and all proofs of the theoretical results are deferred to the Supplement.
2 An ABC-MCMC with global and local proposals
2.1 A review of ABC-MCMC
@@
We consider a given observed data set $\boldsymbol{y}\in\mathbb{Y}$, which is realized from a parametric model $\left\{P_{{\boldsymbol{\theta}}_{0}}:{\boldsymbol{\theta}}_{0}\in\right.$ $\left.\Theta\subset\mathbb{R}^{p}\right\}$. Assume that for each ${\boldsymbol{\theta}}\in\Theta$, $P_{\boldsymbol{\theta}}$ admits a density $p(\cdot\mid{\boldsymbol{\theta}})$ that cannot be directly evaluated but can be sampled from. The ABC posterior can be achieved by augmenting the target posterior from $\pi_{\varepsilon}({\boldsymbol{\theta}}\mid\boldsymbol{y})$ to

### v2

# An adaptive approximate Bayesian computation MCMC with Global-Local proposals

@@ Abstract
In this paper, we address the challenge of Markov Chain Monte Carlo (MCMC) algorithms within the approximate Bayesian Computation (ABC) framework, which often get trapped in local optima due to their inherent local exploration mechanism. We propose a novel Global-Local ABC-MCMC algorithm that combines the “exploration” capabilities of global proposals with the “exploitation” finesse of local proposals. We integrate iterated importance resampling into the likelihood-free framework to establish an effective global proposal distribution. For high-dimensional parameter spaces, we optimize the efficiency of the local sampler by leveraging Langevin dynamics and common random numbers. Furthermore, we introduce two adaptive schemes to enhance the algorithmic performance. The first scheme divides the update target of the importance proposal into a sequence of intermediate target distributions that progressively approximate the ABC posterior, thereby gradually updating the importance proposal distribution during the iterations. The second adaptive scheme automatically selects the optimal mixture of global and local moves through sequential optimization, based on a relative version of the expected squared jumping distance (ESJD). We theoretically and numerically demonstrate that our method is able to improve sampling efficiency and achieve more reliable convergence for complex posteriors. We develop a software package that is available at https://github.com/caofff/GL-ABC-MCMC.
Keywords: Approximate Bayesian computation, Metropolis-adjusted Langevin, iterated sampling importance resampling, adaptive proposal, sequential optimization
@@ 1 Introduction
Traditional Bayesian inference typically relies on the assumption that the likelihood functions of statistical models are tractable. However, for many complex applications, the likelihood functions have no explicit expressions or are difficult to estimate numerically. Approximate Bayesian computation (ABC) (Pritchard1999Population; Beaumont2002ABC) is a likelihood-free inference method based on simulations, it only requires to simulate synthetic data from the model without evaluating the likelihood function. Rejection ABC (tavare1997inferring; Pritchard1999Population) is the simplest and most intuitive version of ABC algorithms, which repeatedly draws samples of parameter $\boldsymbol{\theta}$} independently from prior $\pi$, simulates synthetic data $\boldsymbol{x}$ for each sample of $\boldsymbol{\theta}$}, and rejects the parameter $\boldsymbol{\theta}$} if the discrepancy $\mathfrak{D}(\boldsymbol{x},\boldsymbol{y})$ between the observed data $\boldsymbol{y}$ and the simulated data $\boldsymbol{x}$ exceeds a pre-specified tolerance threshold $\varepsilon$. In practice, it is inefficient to sample parameters from a prior distribution, especially for high dimensional cases. Most of samples drawn from the prior distribution would fall into the low posterior region, which leads to very high rejection rates. Then accepting one sample may require thousands or even millions of draws.
@@
Marjoram2003MCMC introduce an ABC Markov chain Monte Carlo (MCMC) algorithm to approximate the posterior distribution. Proposal distributions are typically chosen to generate local moves that depend on the last state of the chain in order to guarantee an admissible acceptance rate. Theoretically, a small tolerance value $\varepsilon$ leads to good posterior approximation. However, designing an efficient proposal to explore the parameter space, particularly in the case of high dimensional multimodal posterior distributions, poses a significant challenge. Wegmann2009MCMC enhanced the performance of ABC-MCMC by relaxing the tolerance, incorporating subsampling, and applying regression adjustments to the MCMC output. Sisson2007SMC; del2012adaptive integrated ABC-MCMC algorithms into the SMC framework (del2006sequential). The sequence of intermediate target distributions is defined by ABC posteriors, with a series of tolerance parameters. del2012adaptive adaptively determines this sequence by managing the divergence of particles. Clarte2020Gibbs; Likelihood-freeapproximateGibbssampling propose Gibbs version of the approximate Bayesian computation approaches that focuses on lower-dimensional conditional distributions. wang2022approximate incorporates a flexible machine learning classifier into ABC, avoiding the need for summary statistics, to enable automated comparison of real and simulated data. Hamiltonian Monte Carlo (HMC) (DUANE1987HMC; neal2011mcmc) employ Hamiltonian dynamics to propose candidate parameters, efficiently avoiding the random walk behavior that can impede the mixing of proposals in high-dimensional spaces. meeds2015HABC explored the application of HMC to ABC and proposed using completely random numbers (CRNs) to improve the stability of gradient estimates to improve Hamiltonian dynamic.
@@
However, for a high dimensional multimodal posterior distribution, all local samplers mix poorly because the exploration is inherently slow and switching between isolated modes can be extremely rare. Independent proposals can generate more global updates to transfer particles between modes, but they are difficult to design, but a poor independent sampler in MCMC means an inefficient random walk, which will cause the algorithm to converge slowly. For likelihood-based Bayesian inference, samsonov2022local-global; gabrie2022adaptive combine the local and global proposals to improve the efficiency of MCMC. cabezas2024markovian introduces an adaptive tempering mechanism in continuous normalizing flows to learn the complex target distribution. Additionally, normalizing flows (rezende2015variational; papamakarios2021normalizing) are used to improve the exploration capabilities of MCMC independent proposals.
In this paper, we propose a novel approach that combines the advantages of global and local MCMC proposals for challenging posterior distributions within a likelihood-free inference framework. Our work is inspired by samsonov2022local-global, but beyond the fundamental differences between likelihood-based inference and approximate Bayesian computation, we develop an alternative structural design for integrating global and local moves. Concretely, at each iteration, the algorithm probabilistically selects either a global or a local proposal such that the particle update is exclusively global or local. Moreover, we propose a sequential optimization approach based on a unit cost version of expected square jump distance (cESJD) criterion, which is then used to adaptively adjust the probability of choosing the global proposal at each iteration. This adaptive strategy enables a flexible combination of global and local kernels based on their relative performance, thereby ensuring improved overall algorithmic efficiency.
Specifically, we propose a likelihood-free iterated-sampling importance resampling (i-SIR) approach to construct an efficient global ABC-MCMC move. By extending the results of samsonov2022local-global, we show that ABC-i-SIR algorithm can be interpreted as a systematic-scan two-stage Gibbs sampler, and prove the uniform geometric ergodicity of the ABC-i-SIR Markov kernel. Furthermore, we also prove the V-geometric ergodicity of the combination of ABC-i-SIR Markov kernel and local Markov kernel, and hence show that under some conditions, the mixing rate of the Global-Local Markov Kernel is better than that of the local Markov kernel.
To further enhance the efficiency of the global ABC-MCMC move, we introduces intermediate target distributions with a decreasing sequence of tolerance within a single Markov chain, progressively and stably improve the proposal distribution using the weighted samples generated by ABC-i-SIR. We adaptively reduce the tolerance thresholds by controlling the proportion of alive training samples. This makes the selection of the proposal distribution more robust.
For high-dimensional parameter spaces, we employ the Metropolis-adjusted Langevin algorithm (MALA) to construct local ABC-MCMC moves. A key challenge in ABC-MALA is the estimation of the log-likelihood gradient, as the likelihood function is intractable. We utilize common random numbers to improve the accuracy and stability of the gradient estimates.
@@
The rest of paper is organized as follows. Section 2 presents the framework of our Global-Local ABC-MCMC, the efficient global MCMC move ABC-i-SIR and the efficient local MCMC move ABC-MALA. Section 3 introduces two adaptive schemes for improving algorithmic efficiency: the adaptive updating of importance proposals and the sequential optimization of hyperparameters. The theoretical properties are established in Section 4. Sections 5 and 6 demonstrate the effectiveness of our method through synthetic examples studies and two real model analysis. Section 7 provides the conclusion, and all proofs of the theoretical results are deferred to the Appendix.
2 An ABC-MCMC with global and local moves
@@
We consider a given observed data set $\boldsymbol{y}\in\mathbb{Y}$, which is realized from a parametric model $\left\{P_{{\boldsymbol{\theta}}_{0}}:{\boldsymbol{\theta}}_{0}\in\right$. $\left.\Theta\subset\mathbb{R}^{p}\right\}$. Assume that for each ${\boldsymbol{\theta}}\in\Theta$, $P_{\boldsymbol{\theta}}$ admits a density $p(\cdot\mid{\boldsymbol{\theta}})$ that cannot be directly evaluated but can be sampled from. The ABC posterior can be achieved by augmenting the target posterior from $\pi_{\varepsilon}({\boldsymbol{\theta}}\mid\boldsymbol{y})\propto p(\boldsymbol{y}\mid{\boldsymbol{\theta}})\pi({\boldsymbol{\theta}})$ to
$\pi_{\varepsilon}({\boldsymbol{\theta}},\boldsymbol{x}\mid\boldsymbol{y})=\frac{\pi(\boldsymbol{\theta})p(\boldsymbol{x}\mid{\boldsymbol{\theta}})K_{\varepsilon}(\boldsymbol{x},\boldsymbol{y})}{\iint\pi(\boldsymbol{\theta})p(\boldsymbol{x}\mid{\boldsymbol{\theta}})K_{\varepsilon}(\boldsymbol{x},\boldsymbol{y})d\boldsymbol{x}d{\boldsymbol{\theta}}},$
where $\boldsymbol{x}$ denotes synthetic data simulated from $p(\cdot\mid{\boldsymbol{\theta}})$, $K_{\varepsilon}(\cdot,\cdot)$ is a kernel function and the threshold $\varepsilon$ plays a role of bandwidth. The marginal posterior distribution $\pi_{\varepsilon}({\boldsymbol{\theta}}\mid\boldsymbol{y})$ can be evaluated by integrating out the auxiliary data set $\boldsymbol{x}$ from $\pi_{\varepsilon}({\boldsymbol{\theta}},\boldsymbol{x}\mid\boldsymbol{y})$. The choice of threshold $\varepsilon$ is a trade-off between accuracy and computational speed. A smaller value of threshold can lead to more accurate ABC posterior, but with higher computational cost. When the threshold value $\varepsilon$ approaches to 0, $p_{\varepsilon}(\boldsymbol{y}\mid{\boldsymbol{\theta}})$ converges to the exact Bayesian posterior $\pi(\boldsymbol{y}\mid{\boldsymbol{\theta}})$.
2.1 An ABC-MCMC with global and local moves
Standard ABC methods based on rejection sampling become infeasible in high-dimensional parameter spaces due to poor acceptance rates. ABC-MCMC offers a more efficient alternative by constructing a Markov chain targeting $\pi_{\varepsilon}(\boldsymbol{\theta},\boldsymbol{x}\mid\boldsymbol{y})$. While local proposals $q(\cdot\mid\boldsymbol{\theta})$ in ABC-MCMC facilitate efficient exploration near the current state, they often suffer from poor mixing for complex or multimodal posteriors, as the chain can become trapped in local modes.
To enhance the exploration capabilities of MCMC and facilitate more frequent transitions between distinct modes, it is necessary to incorporate more global updates. However, designing efficient global proposals is challenging due to the unknown posterior distribution, which is analogous to the difficulty of performing posterior inference itself. Furthermore, proposing each parameter globally and independently often fails to capture the dependencies and correlations among parameters, hindering effective exploration of the parameter space. To address these challenges, we propose a Global-Local ABC-MCMC framework that integrates both local and global proposals. This approach leverages the complementary advantages of each proposal type, striking an optimal balance that enables thorough exploration of local regions while simultaneously ensuring a more comprehensive coverage of the entire parameter space. Specifically, at each iteration, the algorithm alternates between a global move with probability $\gamma$ and a local move with probability $1-\gamma$.
The most common and straightforward local move in ABC-MCMC involves generating a candidate parameter $\boldsymbol{\theta}^{*}$ from a proposal distribution $q(\cdot\mid\boldsymbol{\theta}_{t})$, followed by simulating synthetic data $\mathbf{x}^{*}$ from the model $P_{\boldsymbol{\theta}^{*}}$. The candidate is then accepted with the Metropolis–Hastings acceptance probability
$\alpha\big[(\boldsymbol{\theta},\mathbf{x}),(\boldsymbol{\theta}^{*},\mathbf{x}^{*})\big]=\min\left\{1,\frac{\pi(\boldsymbol{\theta}^{*})K_{\varepsilon}(\mathbf{x}^{*},\mathbf{y})q(\boldsymbol{\theta}^{*}\mid\boldsymbol{\theta})}{\pi(\boldsymbol{\theta})K_{\varepsilon}(\mathbf{x},\mathbf{y})q(\boldsymbol{\theta}\mid\boldsymbol{\theta}^{*})}\right\},$
otherwise the chain remains at the current state. Typically, the proposal distribution $q(\cdot\mid\boldsymbol{\theta}_{t})$ is taken to be a Gaussian distribution centered at the current state $\boldsymbol{\theta}_{t}$. Conversely, if the proposal distribution is independent of the current state $\boldsymbol{\theta}_{t}$, the resulting move corresponds to a global ABC-MCMC update.
Formally, the transition kernel of the Global-Local ABC-MCMC is given by
$k\big((\boldsymbol{\theta},\boldsymbol{x}),(\boldsymbol{\theta}^{*},\boldsymbol{x}^{*})\big)=\gamma\,k_{g}\big((\boldsymbol{\theta},\boldsymbol{x}),(\boldsymbol{\theta}^{*},\boldsymbol{x}^{*})\big)+(1-\gamma)\,k_{l}\big((\boldsymbol{\theta},\boldsymbol{x}),(\boldsymbol{\theta}^{*},\boldsymbol{x}^{*})\big),$
where $k_{l}$ denotes a local ABC-MCMC kernel, and $k_{g}$ is a global kernel that proposes more distant candidates to facilitate large jumps across the parameter space. The parameter $\gamma\in[0,1]$ controls the balance between exploitation (local moves) and exploration (global moves). When $\gamma=0$, the algorithm reduces to the standard ABC-MCMC, relying exclusively on local proposals; when $\gamma=1$, only global proposals are used. By integrating these two complementary mechanisms, the Global-Local ABC-MCMC achieves improved mixing and robustness, enabling efficient sampling from challenging, potentially multimodal approximate posteriors. Algorithm 2.1 illustrates one iteration of the Global-Local ABC-MCMC.
Algorithm 1 Single stage of Global-Local ABC-MCMC
Input: Previous state $({\boldsymbol{\theta}}_{t},\boldsymbol{x}_{t})$, global ABC-MCMC move $k_{g}\big((\boldsymbol{\theta}_{t},\boldsymbol{x}_{t}),\cdot\big)$, local ABC-MCMC move $k_{l}\big((\boldsymbol{\theta}_{t},\boldsymbol{x}_{t}),\cdot\big)$.
1: Sample $u\sim\mathcal{U}[0,1]$.
2:if $u<\gamma$ then
3:  Obtain $({\boldsymbol{\theta}}_{t},\boldsymbol{x}_{t})$ by running a global ABC-MCMC move $k_{g}\big((\boldsymbol{\theta}_{t},\boldsymbol{x}_{t}),\cdot\big)$.
4:else
5:  Obtain $({\boldsymbol{\theta}}_{t},\boldsymbol{x}_{t})$ by running a local ABC-MCMC move $k_{l}\big((\boldsymbol{\theta}_{t},\boldsymbol{x}_{t}),\cdot\big)$.
6:end if
2.2 The global move based on iterated-sampling importance resample
Directly using a proposal distribution $q(\cdot)$ that is independent of the current state to construct global ABC-MCMC moves can indeed facilitate transitions between modes. However, when such an independent proposal diverges significantly from the ABC posterior, the Metropolis–Hastings acceptance probability $\alpha\left[({\boldsymbol{\theta}},\boldsymbol{x}),({\boldsymbol{\theta}}^{*},\boldsymbol{x}^{*})\right]=\operatorname{min}\left\{1,\frac{\pi({\boldsymbol{\theta}}^{*})K_{\varepsilon}(\boldsymbol{x}^{*},\boldsymbol{y})q({\boldsymbol{\theta}})}{\pi({\boldsymbol{\theta}})K_{\varepsilon}(\boldsymbol{x},\boldsymbol{y})q({\boldsymbol{\theta}}^{*})}\right\}$ tends to be very low. This limitation poses challenges for the global exploration capabilities of Global-Local ABC-MCMC, especially in high-dimensional parameter spaces. Therefore, it is essential that the global proposal closely approximates the ABC posterior; however, inferring such a proposal is nearly as complex as inferring the ABC posterior itself.
To address this challenge, we propose an effective global ABC-MCMC move constructed via ABC Sampling Importance Resampling (SIR). SIR is a commonly used technique to transition from a simple distribution to a more complex target distribution. Specifically, the ABC-SIR procedure involves drawing samples ${\boldsymbol{\theta}}_{i}$ from a simple and easy-to-sample proposal distribution $q({\boldsymbol{\theta}})$, and for each ${\boldsymbol{\theta}}_{i}$, generating corresponding simulated data $\boldsymbol{x}_{i}\sim P_{{\boldsymbol{\theta}}_{i}}$, where $i=1,\ldots,N$. Each of these $N$ samples is then assigned an importance weight
$W_{i}\propto\frac{\pi({\boldsymbol{\theta}}_{i})K_{\varepsilon}(\boldsymbol{x}_{i},\boldsymbol{y})}{q({\boldsymbol{\theta}}_{i})}.$
Finally, resampling from these weighted samples yields a set of samples approximately drawn from the target ABC posterior distribution $\pi_{\varepsilon}({\boldsymbol{\theta}}\mid\boldsymbol{y})$.
Here we utilize iterated SIR (i-SIR) to design efficient global ABC-MCMC move. The likelihood-based i-SIR is originally proposed in tjelmeland2004using and further studied in andrieu2010particle; andrieu2018uniform; samsonov2022local-global. Here, we introduce a likelihood-free version of i-SIR as a more efficient global ABC-MCMC moves, refer to as ABC-i-SIR. Algorithm 2 illustrates a single stage of ABC-i-SIR. In one stage of ABC-i-SIR, we sample $\left\{{\boldsymbol{\theta}}_{(j)}^{*}\right\}_{j=1:N_{b}}$ from the independent proposal distribution, and generate corresponding synthetic data $\left\{\boldsymbol{x}_{(j)}^{*}\right\}_{j=1:N_{b}}$. We then combine the newly generated particles with the last state of the chain $({\boldsymbol{\theta}}_{(0)}^{*},\boldsymbol{x}_{(0)}^{*})=({\boldsymbol{\theta}}_{t},\boldsymbol{x}_{t})$ as candidate particles. The new state $({\boldsymbol{\theta}}_{t+1},\boldsymbol{x}_{t+1})$ is obtained by sampling from the candidate particles $\left\{{\boldsymbol{\theta}}_{(j)}^{*},\boldsymbol{x}_{(j)}^{*}\right\}_{j=0:N_{b}}$ with importance weights $w_{j}\propto\pi({\boldsymbol{\theta}}_{(j)}^{*})K_{\varepsilon}(\boldsymbol{x}_{(j)}^{*},\boldsymbol{y})/q({\boldsymbol{\theta}}_{(j)}^{*})$.
Algorithm 2 Single stage of ABC-i-SIR algorithm with independent proposals